OpenStudy (anonymous):

For which values of k are the vectors u = (1, k, 2), v = (-3,4, I), w = (k, 0,l) linearly independent?

OpenStudy (amistre64):

what does linearly independant mean?

OpenStudy (amistre64):

it has something to do with column vectors and a determinant of zero

OpenStudy (anonymous):

linearly independent means the sum of the vectors does *not* equal zero

OpenStudy (amistre64):

\begin{array}c u&v&w\\ 1&-3&k\\ k&4&0\\ 2&l&l\\ \end{array} 1(4.l - l.0) + -3(0.2 - k.l) + k(k.l - 2.4) = N ; if N = 0 they are dependant

OpenStudy (amistre64):

the rest of it id have to reread alot of stuff to remember :)

OpenStudy (anonymous):

cheers bbz

OpenStudy (anonymous):

its weird that you have 2 variables floating around in there =/ another way to show linear independence is to row reduce the matrix to the identity matrix. thats a little rough with the two variables though.

OpenStudy (amistre64):

i wonder if we can trial and error it :)

OpenStudy (anonymous):

joemath, the l's are meant to be number one's

OpenStudy (anonymous):

oh <.< then its not that hard. let me see if i can do it on paper, one sec lol

OpenStudy (amistre64):

.... lol, you know how many synapses i just blew out trying to solve for l and k ...

OpenStudy (anonymous):

and amistre's method might be more efficient. you find the det of the matrix in terms of k, and you dont want it to be zero. so find out what values make it 0, and just say k can be anything but those values.

OpenStudy (amistre64):

\[1(4.1 - 1.0) + -3(0.2 - k.1) + k(k.1 - 2.4) = N\] \[4 + 3k + k^2 - 8k = N\] \[k^2 - 5k+4 = N\]

OpenStudy (anonymous):

just from observation, k = 4 is a bad value.

OpenStudy (anonymous):

thank you all.

OpenStudy (amistre64):

f(k) = 0 when k = 4 or 1